Find an equation of the sphere that has endpoints of a diameter at A and B


A(1,4,-2) B(-7,1,2)


1
Expert's answer
2021-12-13T14:53:56-0500

The centre of the sphere is the midpoint of the diameter


"x_C=\dfrac{x_A+x_B}{2}=\dfrac{1-7}{2}=-3"

"y_C=\dfrac{y_A+y_B}{2}=\dfrac{4+1}{2}=2.5"

"z_C=\dfrac{z_A+z_B}{2}=\dfrac{-2+2}{2}=0"

The center of the sphere is "C=(-3, 2.5, 0)."

The length of the diameter is 


"d=2r=\sqrt{(-7-1)^2+(1-4)^2+(2-(-2))^2}"

"=\sqrt{89}"

So the radius of the sphere is "\sqrt{89}/2."

The equation of the sphere is


"(x+3)^2+(y-2.5)^2+z^2=22.25"